OpenStudy (anonymous):

Convert to base 6: \[3BA.25_{14}\]

OpenStudy (anonymous):

satellite i really need your help on my question...

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

\[3BA.25 = 03BA.25 = 03 B A .2 5 = 03 15 14. 02 05

ganeshie8 (ganeshie8):

\[\left(3BA.25\right)_{14} = 14^2\times 3 + 14\times 11+10 + \frac{2}{14} + \frac{5} {14^2} = \left(752.168\right)_{10}\]

ganeshie8 (ganeshie8):

divide successively by 6 to convert to base 6

OpenStudy (anonymous):

yeah i'm trying to find another way to do it faster

OpenStudy (anonymous):

but you can't used a fixed table for it =/

OpenStudy (anonymous):

@ganeshie8 ... got a question my teacher uses fixed tables when he converts base 3 to base 9 or base 9 to base 3 ... why can you not use it in this instance

ganeshie8 (ganeshie8):

I don't know what you mean by fixed table but converting back and forth between base b and base b^n is trivial, you just need to fix the number of positions. conversion between any other bases is not trivial, you need to divide. for this particular problem 14 is not a power of 6, so you don't have a choice, you need to divide by 6 and collect remainders.

OpenStudy (anonymous):

alright makes sense